a man cycles for sometime at 16 km/h and returns at 15 km/h. the total time taken is 31/4 hours. find the distance he had cycled together.
Answer by Alan3354(69443)

a man cycles for sometime at 16 km/h and returns at 15 km/h. the total time taken is 31/4 hours. find the distance he had cycled together.
-------------
Avg speed for a round trip is
2*16*15/(16+15) = 480/31 km/hr
RT distance = (480/31)*(31/4) = 120 km